1. How many gallons of a 10% solution. Would need to be added to a 20% solution to obtain 50 gallons of a 30% solution?
2.How many gallons of milk containing 4% fat is mixed with milk containg 1% fat to obtain a 15 gallon mixture containing 2% fat?

Answer by Alan3354(69443)

1. How many gallons of a 10% solution. Would need to be added to a 20% solution to obtain 50 gallons of a 30% solution?
It's not possible. Adding 10% to 20% will reduce the concentration.
Any mix of 2 concentrations will always result in a %age between the two.
-------------------
2.How many gallons of milk containing 4% fat is mixed with milk containg 1% fat to obtain a 15 gallon mixture containing 2% fat?
0.04*x + 0.01*(15-x) = 0.02*15
4x + 15-x = 30
x = 5 gallons of 4%
10 gallons of 1%